OpenStudy (anonymous):

‎1:52, 2:33, 4:31, 5:12, ?

OpenStudy (anonymous):

honestly sathi, I'm too sleepy to think lol :)

OpenStudy (anonymous):

this is last question b4 i go to bed

OpenStudy (anonymous):

the only thing I got is, 52 - 33 = 19, 31-19 = 12 etc

OpenStudy (anonymous):

can it be 6:0?

OpenStudy (anonymous):

sicne you keep on subtracting the previous terms

OpenStudy (anonymous):

since*

OpenStudy (anonymous):

1:52 2:33 (52-33) = 19 4:31 (31-19) = 12 5:12 (12-12) __: 0 something and 0? I'm not sure

OpenStudy (anonymous):

makes any sense?

OpenStudy (anonymous):

na

OpenStudy (anonymous):

>_>

OpenStudy (anonymous):

atleast I've tried lol

OpenStudy (anonymous):

I'M SLEEPY! e_e

OpenStudy (anonymous):

thats good

OpenStudy (anonymous):

u r a deserving student

OpenStudy (anonymous):

lol my pc is toooooo slow..

OpenStudy (anonymous):

I am a deserving student? .-.

OpenStudy (anonymous):

anwar! =-=

OpenStudy (anonymous):

yeah?

OpenStudy (anonymous):

I said, I'm not mad at you

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

hey guys share answer yar

OpenStudy (anonymous):

and I'm sorry about earlier, didn't have to overreact

OpenStudy (anonymous):

It's ok.

OpenStudy (anonymous):

and you must not take your jokes too far

OpenStudy (anonymous):

>_>....lol :)

OpenStudy (anonymous):

Hay!! :)

OpenStudy (anonymous):

=P :)

OpenStudy (anonymous):

hwy guyws wtf u r talkin g m waitng for answer lol

OpenStudy (anonymous):

oh nothin'

OpenStudy (anonymous):

I seriously have no answer lol

OpenStudy (anonymous):

and I seriously want to go to bed

OpenStudy (anonymous):

and I'm seriously sorry

OpenStudy (anonymous):

me too lol...

OpenStudy (anonymous):

good night

OpenStudy (anonymous):

Me three!!

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

lol, good night ^_^

OpenStudy (anonymous):

go to bed and check your question tomorrow sathi :)

OpenStudy (anonymous):

tomorrwow mng is my exam lol

OpenStudy (anonymous):

well get some sleep, forget abt it

OpenStudy (anonymous):

i have to score highest mark in that exam lol

OpenStudy (anonymous):

you will, if you get some sleep lol

OpenStudy (anonymous):

ok u slee[ i will have some more tasks and than join u lol.

OpenStudy (anonymous):

join me? =-=

OpenStudy (anonymous):

OMG!!

OpenStudy (anonymous):

I'll just leave lol

OpenStudy (anonymous):

idr

OpenStudy (anonymous):

ok bye guyes///

OpenStudy (anonymous):

uzuma will try to help you out~ ciao now ~

OpenStudy (anonymous):

ciao tomooorow lol

OpenStudy (anonymous):

it might b 6:11 :)

OpenStudy (anonymous):

but dont know it makes sense or not :S

OpenStudy (anonymous):

I thought about this option, but yeah there are not enough evidence for it. the numbers add up to 8. But what about 7:10?

OpenStudy (anonymous):

yea right there must b some more logic too

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

ok 1/52,2/33,4/31,5/12 and 1+1=2, 2+2= 4, 4+1=5 so the pattern is 1,2,1,2.... and 52 -19=33,33-2=31,31-19=12 so the pattern is -19,-2,-19,-2..... therefore next is 7/10, 8/-9.....